Can some one show me how to do this problem
using the five-step problem
For a holiday party, Marcus is making a five-layer torte, which consists of five 9-inch diameter cake layers with a cream topping on each layer. He knows that one recipe of cream will cover 60 square inches of cake. How many batches of the topping must he make to complete his torte? (Use п = 3.14 as an approximate value) im lost please

Answer by mananth(16949)

Area of a circle = pi*r^2
r=4.5 in
Area of 1 layer torte = 3.14*4.5 ^2
Area of 1 layer = 63.58in^2
area of 5 layers = 5*63.58 =317.92 in ^2
..
60sq. in --- 1 batch
317.92 sq. in you will require 317.92/60
=5.92 batches
6 batches must be taken